OpenStudy (anonymous):

y = 2x^3 - 4x^2 I just need this for clarification. Solve for x intercept

OpenStudy (anonymous):

in order to solve for y-intercept, set the y to 0 0=2x^3-4x^2 0=2x^2(x-2) 2x^2=0 so x=0 x-2=0 so x=2 the x-intercepts of the function are 0 and 2

OpenStudy (anonymous):

Ok that's what I got, thank you

OpenStudy (anonymous):

no problem :D